You are working in a virology lab and are given a sample of high-titer influenza type A virus and Madin-Darby canine kidney (MDCK) epithelial cells which you seed into tissue culture plates. After culturing the MDCK cells to obtain confluent monolayers, your task is to determine the virus titer and express this titer in pfu/1000uL. You therefore make four 100-fold dilutions and one 10-fold dilution of the neat virus stock and plate 500uL of the final dilution. After 72 hours, you remove the agar overlay, fix and then stain the infected cells with crystal violet. The plates corresponding to your final dilution above contain an average of 27 plaques. What is the titer of the original virus stock?

(A) 5.4 × 1010

(B) 2.7 × 1010

(C) 13.5 × 108

(D) 5.4 × 107

(E) 2.7 × 106
Answer #1

Titer of the original virus stock = Number of plaques / 1 / Dilution factor

Dilution factor = 1/100 X 1/100 X 1/100 X 1/100 X 1/10 = 1/10^9

Plate dilution factor = 500/1000 = 1/2

Titer of the original virus stock = 27 / 1 / 1/10^9 X 1/2

                                            = 54 X 10^9

                                            = 5.4 X 10^10